# TRIDENT: Tri-Modal Molecular Representation Learning with Taxonomic Annotations and Local Correspondence

### v1

@@ 13 Classification path retrieval
14 Endpoint {PubChem /classification_2.fcgi?hid={hid}&search_uid={cid}}
15 Path construction {recursive traversal of parent-child nodes}
16 Output {hierarchical path string, description}
17 Result processing
18 Data structure {CID → {classification_system → {path, description}}}
19 Intermediate saving {save after each batch for resumability}
20 Error handling {log warnings and errors, continue processing}
21 Output
22 JSON file {complete taxonomy annotation for all CIDs}
@@ Appendix B Experimental Setup
3 SMILES encoder {MoLFormer (768-dim)}
4 Text description encoder {SciBERT (768-dim)}
5 Category encoder {shared with text encoder (768-dim)}
6 Projection layers
7 SMILES projection {Linear→GELU→LayerNorm→Dropout→Linear→GELU→LayerNorm→Linear (512-dim)}
8 Text projection {Linear→GELU→LayerNorm→Dropout→Linear→GELU→LayerNorm→Linear (512-dim)}
9 Global contrastive loss (GRAM3Modal)
10 Volume computation {determinant of Gram matrix}
@@
12 Volume-based alignment {cross entropy on negative volumes}
13 InfoNCE alignment {standard contrastive across modalities}
14 Local functional group alignment
15 Functional group detection {RDKit Fragments}
16 FG representation {weighted pooling of fragment embeddings}
17 FG contrastive loss {local InfoNCE between SMILES and text}
18 Momentum-based loss integration
@@ 21 Dynamic update $\alpha=\beta\cdot\alpha_{prev}+(1-\beta)\cdot(global\_loss/total\_loss)$}
22 Training configuration
23 Optimization {Adam (lr=1e-5)}
24 Encoder freezing {both text and SMILES encoders}
25 Distributed training {DDP, NCCL backend}
26 Batch size {40 per GPU, multi-GPU}
@@ B.1 Model Architecture
    • Global contrast: Applies GRAM3Modal method to calculate volume loss and InfoNCE loss across three modalities
@@ C.2 TDC Datasets
TDC (Therapeutics Data Commons) is a large-scale dataset collection specifically designed for therapeutics research, providing more challenging and practically valuable molecular property prediction tasks. Table 7 presents the detailed information of the 5 TDC datasets we selected.
@@ Appendix E Additional Results
In this section, we present additional experimental results that complement the main findings reported in the paper. These include performance evaluations on larger-scale datasets from the TDC benchmark, more extensive ablation experiments, and additional analyses that provide deeper insights into TRIDENT’s capabilities.
E.1 Performance on Larger TDC Datasets
While the main paper focused on smaller TDC datasets to demonstrate TRIDENT’s data efficiency, we also evaluated our method on larger-scale molecular property prediction tasks. Table 8 presents the results on the AMES mutagenicity dataset (7,255 molecules) and the hERG cardiotoxicity dataset (648 molecules). In summary, TRIDENT’s superior performance on both the large-scale AMES dataset and the moderately-sized hERG dataset demonstrates the versatility and scalability of our approach. The consistent improvements across different dataset sizes—from hundreds to thousands of molecules—validate that the tri-modal alignment strategy and hierarchical taxonomic annotations provide robust molecular representations that generalize well across various scales and prediction tasks. These results complement our findings on larger datasets and further establish TRIDENT as a powerful framework for molecular property prediction across the full spectrum of practical applications in drug discovery.
@@ E.1 Performance on Larger TDC Datasets
TRIDENT (M-S) 85.37±0.30 78.74±0.50 87.60±1.20 81.11±2.64
TRIDENT (M-M) 86.87±0.60 80.20±1.44 83.31±1.63 83.33±2.62
E.2 Performance without LLM Summary
@@
E.3 Impact of Tri-modal vs. Concatenated Text Architecture

### v2

@@ 15 Classification path retrieval
16 Endpoint .{PubChem /classification_2.fcgi?hid={hid}&search_uid={cid}}
17 Path construction .{recursive traversal of parent-child nodes}
18 Output .{hierarchical path string, description}
19
201exResult processing
21 Data structure .{CID → {classification_system → {path, description}}}
22 Intermediate saving .{save after each batch for resumability}
23 Error handling .{log warnings and errors, continue processing}
24
251exOutput
26 JSON file .{complete taxonomy annotation for all CIDs}
@@ Appendix B Experimental Setup
3 SMILES encoder .{MoLFormer (768-dim)}
4 Text description encoder .{SciBERT (768-dim)}
5 Category encoder .{shared with text encoder (768-dim)}
6
71exProjection layers
8 SMILES projection .{Linear→GELU→LayerNorm→Dropout→Linear→GELU→LayerNorm→Linear (512-dim)}
9 Text projection .{Linear→GELU→LayerNorm→Dropout→Linear→GELU→LayerNorm→Linear (512-dim)}
10
111exGlobal contrastive loss (GRAM3Modal)
12 Volume computation .{determinant of Gram matrix}
@@
14 Volume-based alignment .{cross entropy on negative volumes}
15 InfoNCE alignment .{standard contrastive across modalities}
16
171exLocal functional group alignment
18 Functional group detection .{RDKit Fragments}
19 FG representation .{weighted pooling of fragment embeddings}
20 FG contrastive loss .{local InfoNCE between SMILES and text}
21
221exMomentum-based loss integration
@@ 25 Dynamic update $\alpha=\beta\cdot\alpha_{prev}+(1-\beta)\cdot(global\_loss/total\_loss)$}
26
271exTraining configuration
28 Optimization .{Adam (lr=1e-5)}
29 Encoder freezing .{both text and SMILES encoders}
30 Distributed training .{DDP, NCCL backend}
31 Batch size .{40 per GPU, multi-GPU}
@@ B.1 Model Architecture
    • Global contrast: Applies GRAM3Modal method to calculate volume loss and InfoNCE loss across three modalities.
@@ C.1 MoleculeNet Datasets
ESOL 1,127 Water Solubility Predicts aqueous solubility (log S), fundamental for drug formulation and delivery
FreeSolv 641 Solvation Free Energy Predicts hydration free energy, important for understanding molecular interactions
Lipophilicity 4,200 Lipophilicity Predicts octanol-water partition coefficient (log D), crucial for drug absorption and distribution
@@ C.2 TDC Datasets
TDC (Therapeutics Data Commons) is a large-scale dataset collection specifically designed for therapeutics research, providing more challenging and practically valuable molecular property prediction tasks. Table 7 presents the detailed information of the 7 TDC datasets we selected.
@@
CYP P450 2C19 12,665 Drug Metabolism Prediction Predicts inhibition of CYP2C19 enzyme, essential for assessing drug-drug interactions
Caco-2 906 Permeability Prediction Predicts intestinal permeability using Caco-2 cell line, critical for oral drug bioavailability
@@ Appendix E Additional Results
In this section, we present additional experimental results that complement the main findings reported in the paper. These include performance evaluations on MoleculeNet regression tasks, larger-scale datasets from the TDC benchmark, more extensive ablation experiments, and additional analyses that provide deeper insights into TRIDENT’s capabilities.
E.1 Performance on MoleculeNet Regression Tasks
To further demonstrate TRIDENT’s effectiveness across different task types, we evaluate our method on three regression benchmarks from MoleculeNet: ESOL (water solubility), FreeSolv (solvation free energy), and Lipophilicity (octanol-water partition coefficient). As shown in Table 8, TRIDENT (M-M) achieves the best performance on ESOL and FreeSolv, and matches the state-of-the-art performance on Lipophilicity. These results demonstrate that TRIDENT’s multimodal learning framework is effective not only for classification tasks but also for continuous property prediction.
Table 8: Regression performance (RMSE) on MoleculeNet benchmark. Lower values indicate better performance.The best results are marked in bold, and the second-best are underlined.
Dataset Uni-Mol BioT5 BioT5+ MolXPT MolFormer MolT5 TRIDENT(M-M)
[…]
E.2 Performance on Larger TDC Datasets
While the main paper focused on smaller TDC datasets to demonstrate TRIDENT’s data efficiency, we also evaluated our method on larger-scale molecular property prediction tasks. Table 9 presents the results on the AMES mutagenicity dataset (7,255 molecules) and the hERG cardiotoxicity dataset (648 molecules). Additionally, Table 10 reports performance on the CYP P450 2C19 inhibition dataset (12,665 molecules) and the Caco-2 permeability regression dataset (906 molecules).
In summary, TRIDENT’s superior performance across these diverse large-scale datasets—from the moderately-sized hERG and Caco-2 to the large-scale AMES and CYP P450 2C19, demonstrates the versatility and scalability of our approach. The consistent improvements across different dataset sizes, ranging from hundreds to over ten thousand molecules, and across both classification and regression tasks, validate that the tri-modal alignment strategy and hierarchical taxonomic annotations provide robust molecular representations that generalize well. These results complement our findings on larger datasets and further establish TRIDENT as a powerful framework for molecular property prediction across the full spectrum of practical applications in drug discovery.
@@ E.2 Performance on Larger TDC Datasets
[…]
Table 10: Performance on larger-scale TDC datasets. For CYP P450 2C19, we report accuracy (ACC) and ROC-AUC. For Caco-2 regression task, we report RMSE (lower is better).
Dataset Metric MolT5 MolXPT BioT5 BioT5+ TRIDENT(M-M)
[…]
E.3 Performance without LLM Summary
@@
E.4 Impact of Tri-modal vs. Concatenated Text Architecture